40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9193     Entry time: Thu Oct 3 02:51:26 2013
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI locked "forever", some ALS fiddling 

First up for me this evening was getting the PRMI locked.

I used the IFO configure screen to lock the X and Y arms, then aligned them using the ASS scripts.  Then used the IFO config screen to restore the Michelson, and did some fine tune tweaking of the BS alignment by looking at the AS camera.  Then, I restored the PRMI from the IFO config screen, tweaked the PRM a little bit in yaw, and was able to get a lock using REFL 165 I&Q for ~25 minutes before I got bored and unlocked things.  I used the ASS for the PRM to align the PRM, then turned off the ASS.  POP110 and POP22 both drifted down, but by a small amount, and at the end (when I turned the ASS back on for PRM), they picked back up to about their original levels.

PRMI_lock_ASSatBeginningAndEnd.png

(Note to self: to get it to print both plots, chose custom paper size, make it 14.5 by 11.  Don't ask why, just do it, because it works.  Also, in PNG device properties, increase the compression to 9.)

 

After I played with the PRMI, I started looking at the ALS system. 

I had both arms locked on IR using the regular LSC system (so POX and POY for the error signals).  Then I opened up the green shutters, and got both arms locked on green (so the green lasers were just following the arms...no digital ALS business).  I went out to the PSL table and tweaked up the alignment of the green beams (didn't need much at all, just an itsy bitsy bit in yaw, mostly).  I saw a very strong peak for the Yarm vs. PSL (around -19dBm), and there was a harmonic of that beat.  Opening and closing the Xarm green shutter had no effect on these peaks, so there wasn't any kind of X-Y cross beat sneaking around that I could see.  That's really as far as I got - I think (but haven't checked) that Manasa may have removed the power splitter / combiner, so that the RF analyzer is only looking at the Y beat PD (she mentioned earlier today that she was going to give that a try to narrow things down). 

After that, Rana and I went back to the PRMI for some noise stuff, and worked on the PMC.  See those separate elogs for info on those activites.

ELOG V3.1.3-